OpenStudy (anonymous):

find a_18 when a_1=27, d=-6

OpenStudy (jhannybean):

Using an arithmetic sequence that follows the formula: \(a_n =a_1 +(n-1)d\)

OpenStudy (anonymous):

-75 is my answer

OpenStudy (jhannybean):

\[\large a_{\color{red}{n}} = a_{\color{red}{18}}\] Therefore we can conclude \(n=18\). Let's see..

OpenStudy (jhannybean):

That is correct :)

OpenStudy (anonymous):

thank you, but i did it the long way, d=-6 a_1=27 a_n=?

OpenStudy (jhannybean):

Perhaps you can show me how you've solved it?

OpenStudy (anonymous):

i did 27-6=21 21-6=15 15-6=9 and so on until i got to a18.

OpenStudy (anonymous):

but what is a_n?

OpenStudy (jhannybean):

That's also a good way to go about finding it.

OpenStudy (anonymous):

it cant be 27(18-1)-6.

OpenStudy (jhannybean):

And why do you think it cant be?

OpenStudy (anonymous):

because i got 453

OpenStudy (anonymous):

and -2754 if i do the negative sign, first answer (453) i did a minus sign

OpenStudy (jhannybean):

This is how I solved it.\[a_n = a_{18} \longrightarrow n=18\]\[a_n=27+(18-1)(-6)\]\[a_n=27 -102\]\[a_n=-75\]

OpenStudy (anonymous):

thats why, i forgot the addition sign.

OpenStudy (jhannybean):

:)

OpenStudy (anonymous):

thank you ma'am!

OpenStudy (jhannybean):

No problem and good luck!